When a group of bushwalkers become lost they decide to launch a distress flare vertically into the air from the top of a cliff. at the t seconds after firing, the height h metres, of the flare above the water was h=-5t2+30t+35
(i) What was the height of the flare above the water at t=0 seconds?
(ii) How many seconds after firing did the flare reach its maximum height?
(iii) what was the maximun height when the flare was above:
(a) the water
(b) the top of the cliff

Answer by stanbon(75887)

When a group of bushwalkers become lost they decide to launch a distress flare vertically into the air from the top of a cliff. at the t seconds after firing, the height h metres, of the flare above the water was h(t)=-5t2+30t+35
(i) What was the height of the flare above the water at t=0 seconds?
h(0) = 35 meters
----------------------
(ii) How many seconds after firing did the flare reach its maximum height?
t = -b/2a = -30/(2*-5)= 3 seconds
-----------------
(iii) what was the maximun height when the flare was above:
h(3) = -5*3^2+30*3+35 = -45+90+35 = 80 meters
-------------------
(a) the water
h(3)+35 = 115 meters
-----------------------
(b) the top of the cliff
h(3) = 80 meters
========================
Cheers,
Stan H.